# Nautilus: Locality-aware Autoencoder for Scalable Mesh Generation

### v4

@@ 3.2. Dual-Stream Point Conditioner
Following (Chen et al., 2024b, a, c), our default input condition for mesh generation is point cloud, given its easy accessibility and rich geometric information. In Nautilus, we design a dual-stream point conditioner that provides both global and local geometric guidance.
@@
Global Point Cloud Encoder. In order to capture high-quality global features from the input point cloud, we employ the Michelangelo point cloud encoder (Zhao et al., 2024) to extract complete geometric information. As illustrated in Fig. 5, the point cloud of the entire mesh is encoded into a global latent feature, $f_{\text{glb}}$, which serves as the key and value in the cross-attention layer of the decoder. This global feature encapsulates the overall geometry of the mesh asset, enabling the decoder to grasp the general shape of the input point cloud and autoregressively generate an approximate mesh output that reflects the underlying structure.
@@
Local Point Cloud Encoder. As discussed in Sec. 1, fine-grained local geometric fidelity is crucial for preserving mesh topology and structural coherence. To this end, we propose a local-aware feature extraction mechanism that aligns with our shell-based generation paradigm, providing precise geometric guidance at the vertex-fan level. As depicted in Fig. 5, we employ a PointConv (Zhang et al., 2022) module $f_{\text{loc}}(\cdot)$ to capture local geometric information from the point cloud input. Furthermore, we design a feature injection scheme that augments token embeddings with their corresponding local geometric context, inspired by ControlAR (Li et al., 2024a). Specifically, for each vertex $P_{k,i}$ in the $k$-th shell, we incorporate the local feature $f_{\text{loc}}(O_{k})$ extracted from its central vertex $O_{k}$ into its tokens feature. For each shell’s center vertex, we construct a local neighborhood through KNN sampling of the 100 nearest points, ensuring dense coverage of local surface geometry. This tight integration with our shell-based tokenization enables progressive incorporation of local geometric constraints during generation, enhancing both topological consistency and structural fidelity.
@@
Figure 5. Illustration of our Dual-Stream Point Conditioner. The Global Encoder extracts global features $f_{\text{glb}}$ from the entire point cloud, used as key and value in cross-attention layers. The Local Encoder captures fine-grained topology information via point convolution, adding local geometric features $f_{\text{loc}}$ to each token during generation.
3.3. Autoregressive Sequence Decoder
@@
The core generator of our Nautilus is a transformer decoder that autoregressively predicts the probability of the next mesh token based on the preceding, where each layer consists of a cross-attention layer and a self-attention layer with feed-forward network.
@@
Autoregressive Mesh Token Generation. After tokenization, the mesh generation task is simply reformulated as an autoregressive sequence prediction problem (Vaswani, 2017; Siddiqui et al., 2024; Bai et al., 2024). Using the next-token prediction paradigm, models are trained to predict each discrete coordinate $c\in\{u,v\}$ based on the conditional probability given previously generated tokens. The probability of a mesh $\mathbf{M}$ can be expressed as:
(5) $p(\mathbf{M})=\prod_{i=1}^{L}p(c_{i}|c_{<i}),\quad c_{i}\in\{0,1,\dots,\alpha-1\}$
@@
(6) $L_{\text{CE}}=\text{CrossEntropy}(\hat{S},S(\mathbf{M})_{>0})$
@@
3.4. Image Conditioned Generation
@@
To further expand practical applications, we extend our Nautilus-style tokenization and autoregressive framework to support image-conditioned (Bai et al., 2023) mesh generation. We achieve this by leveraging the Michelangelo (Zhao et al., 2024), a pre-trained shape generation framework that aligns multiple modalities in a unified feature space. Specifically, we replace the original decoder in Michelangelo with our mesh decoder, which incorporates Nautilus-style tokenization. The decoder is trained while keeping the point cloud encoder frozen, enabling mesh generation from the unified feature space. After training, we integrate Michelangelo’s image encoder to enable single-image conditioning. Using the aligned feature space, our decoder generates high-quality meshes directly from images, as demonstrated by the experimental results in Sec. 4.
@@
Figure 6. Qualitative comparison on point cloud conditioned generation between Nautilus and latest open-source methods (Chen et al., 2024b, c). For the last three row, the source dense meshes used to sample the input point cloud are shown in the left-most column.
4. Experiments
4.1. Implementation Details
Our Nautilus framework is trained on a curated dataset of 311K high-quality artist-created mesh assets, which is sourced from our manually cleaned Objaverse data and self-collected data, and each containing up to 8,000 faces, with further statistics are provided in Appendix. For each training asset, we sample 4,096 points as input condition. Nautilus was trained on 8 NVIDIA L40 GPUs for 2 weeks and 400K steps, using Adam optimizer with a learning rate of 0.0001 and a batch size of 16. The transformer decoder includes 24 layers and contains around 500M parameters. During mesh generation, we employ probabilistic sampling with a temperature of 0.5 for stability, generating meshes of 5,000 faces in 4 minutes.
4.2. Qualitative Results
Point Cloud Condition. We compare our Nautilus with the latest open-source direct mesh generation methods (Chen et al., 2024b, c) in the point cloud conditioned generation, where EdgeRunner (Tang et al., 2024) is excluded from this comparison due to its unavailablity of pretrained checkpoints. To construct the test set, we use open-source image-to-3D methods (Team, 2025) to generate dense meshes and directly sample point clouds from them, ensuring that all test samples are entirely unseen during training. The selected samples present significant challenges, featuring thin structures, anisotropic faces, and intricate geometric details.
As shown in Fig.6, existing methods struggle with these challenging cases, exhibiting structural defects such as surface holes and missing components, while failing to capture fine topological details. In contrast, Nautilus generates aesthetically appealing, artist-like meshes that faithfully preserve the geometric details of the input. This performance gap stems from their inadequate locality modeling. Current tokenization approaches generate excessive sequence lengths, limiting these methods to training on simple meshes with fewer than 1,600 faces—insufficient for complex test cases. Furthermore, the limited local dependency compromises structural fidelity and increases manifold defects. Nautilus addresses these limitations by leveraging locality in both mesh tokenization and sequence generation, achieving superior quality even for complex topologies, with additional results presented in Fig.10.
Single-view Image Condition. In Fig. 7, we present a comparison of meshes generated from single-view images between our Nautilus and EdgeRunner (Tang et al., 2024), the only autoregressive method capable of directly conditioning on images  [^2] Note that our comparison is limited to EdgeRunner’s released samples, as their checkpoints are not publicly available.. As shown in Fig. 7, Nautilus generates detailed, manifold meshes with sharp features that accurately match the input conditions, which highlights the superiority of our Nautilus-style tokenization, as further validated through comparisons in Sec. 4.3.
@@ 4.2. Qualitative Results
Figure 7. Qualitative comparison on single image conditioned generation between Nautilus and EdgeRunner (Tang et al., 2024). The results are presented with corresponding normal maps to highlight topological details.
Table 1. Quantitative comparison with other tokenization algorithms. Our Nautilus-style tokenization achieves more efficient compression while better preserving the proximity of neighboring vertices in the sequence.
Metrics AMT (2024c) EdgeRunner (2024) Ours
@@
4.3. Quantitative Results
Tokenization Algorithm. In Table 1, we compare our Nautilus-style mesh tokenization algorithm with Adjacent Mesh Tokenization (AMT) (Chen et al., 2024c) and EdgeRunner (Tang et al., 2024) in two aspects: (1) Compression Ratio measures the reduction in sequence length relative to the original representation $9N$. Higher compression ratios enable modeling of meshes with more faces and better topological complexity. (2) Local Ratio, a new metric that we introduce, quantifies structural coherence among neighboring tokens. For each vertex, we calculate the proportion of its neighboring vertices in the mesh that appear within 100 preceding positions in the sequence. Higher local ratios indicate better preservation of local dependencies after serialization. In Table 1, our tokenization significantly outperforms other methods in both metrics, which is attributed to the explicit locality modeling in its design. In contrast, AMT processes mesh faces in an order of elongated strips rather than interconnected local parts, neglecting local dependencies during tokenization, while EdgeRunner relies on the excessive introduction of specialized tokens, leading to suboptimal performance in both metrics.
Table 2. Quantitative Comparison with latest open-source methods, where our Nautilus framework significantly outperform others in both condition fidelity and visual quality.
@@ 4.3. Quantitative Results
Generation Performance. We compared our point cloud conditioned generation results with MeshAnything (Chen et al., 2024b) and MeshAnythingV2 (Chen et al., 2024c), on a high-quality test set of 500 samples collected using the same protocol as in our qualitative experiments. Following the settings of previous work (Tang et al., 2024), we evaluate overall reconstruction quality by uniformly sampling 1,024 points from the surfaces of both ground truth and predicted meshes, and computing the Chamfer Distance (C.Dist.) and Hausdorff Distance (H.Dist.) between the sampled point clouds. Lower distances between point sets indicate higher reconstruction accuracy and better fidelity to the input condition. In addition, we conducted a user study to assess the proportion of results that users find satisfactory, with further details provided in Appendix. In Table 2, Nautilus significantly outperforms previous approaches across all metrics, validating its effectiveness that is consistent with the qualitative analysis presented in Sec. 4.2.
@@
Figure 8. We conduct ablation comparison on four specific configurations: using only the Coordinate Compression (Comp. only), using only the Nautilus Shell representation (Rep. only), using our Full Tokenization but without the local point encoder, and our Full Version.
4.4. Ablation and Discussion
@@
To validate the effectiveness of our proposed Nautilus-style tokenization and the Dual-Stream Point Conditioner, we perform extensive ablation studies on the tokenization scheme and the local encoder in our conditioner, where an additional quantitative ablation comparison is presented in Appendix.
@@
Tokenization Scheme. We evaluated two configurations to demonstrate the key components of our tokenization design: using only coordinate compression (Comp. only) and using only the Nautilus Shell representation (Rep. only). Fig. 8 compares these variants with our full version. Comp. only achieves a compression ratio of 0.549 and Rep. only achieves 0.484, both substantially higher than our full version’s 0.275. In terms of generation quality, Comp. only fails to construct the topology completely, while Rep. only produces a coarse outline of the target shape. In contrast, with the complete tokenization, our full version produces precisely structured results that are consistent with the input condition in geometry. With the explicit modeling of local neighboring vertices used in our Nautilus Shell representation, both Rep. only and our full version demonstrate better performance compared to Comp. only. In addition, as shown in the comparison among the two ablated groups and our full version, insufficient compression restricts the inclusion of complex training samples, leading to the model’s inability to learn intricate topologies effectively and resulting in generation failures.
@@
Local Point Encoder. To evaluate the local encoder’s effectiveness in the point conditioner, we compare our full version with a variant that uses full tokenization but excludes the local encoder (Full Tokenization). As shown in Fig. 8, the variant without the local encoder exhibits minor manifold defects in complex topological regions, particularly surface holes in local structures. In contrast, our full version, which incorporates the local encoder, successfully resolves these issues and generates precise, compact meshes. The results demonstrate that the local geometric modeling provided by our local encoder yields significant improvements in local details and connectivity, underscoring the importance of enhanced local dependency for generating precise local structures in challenging regions and mitigating manifold defects.
5. Conclusion and Future Works
@@
In this paper, we present Nautilus, a locality-aware autoencoder designed for scalable high-fidelity mesh generation. By leveraging the locality of manifold meshes, Nautilus introduces a novel tokenization algorithm that preserves local dependency while achieving significant sequence compression, as well as a Dual-Stream Point Conditioner that enhances both global consistency and local structure fidelity. Our extensive experiments demonstrate that Nautilus outperforms state-of-the-art methods across various metrics, enabling the generation of meshes with unprecedented topological complexity and details. These results validate the importance of incorporating locality in tokenization and sequence prediction to achieve superior mesh generation. For future work, we aim to extend Nautilus to support broader 3D generation tasks, such as interactive scene editing and multi-object synthesis. Additionally, we plan to explore integrating multimodal conditioning, such as combining image, text, and point cloud inputs, to enhance the versatility of our framework. Another direction is to further optimize the efficiency of Nautilus to facilitate real-time applications.
@@
Figure 9. We release additional point cloud conditioned generation results of our Nautilus across an extensive range of object categories.
@@
Figure 10. We release additional image conditioned generation results of our Nautilus across an extensive range of object categories. The results are presented alongside corresponding normal maps to highlight the topology.
Appendix A Step-by-Step Demonstration of Our Tokenization
@@
1:Input: manifold mesh asset $\mathcal{M}=(\mathcal{V},\mathcal{F})$ consisting of vertices $\mathcal{V}=\{v_{i}\}$ and faces =$\mathcal{F}=\{f_{i}\}$.
@@
11:      Extend $S(\mathcal{M})\leftarrow S(\mathcal{M})\cup\{O_{k},P_{k,1},\dots,P_{k,\text{end}}\}$, completing shell $k$.
@@
14:      If $\deg(N^{\text{max}}_{k,\text{end}})>4$ (to ensure at least three connected faces in the next shell):
@@
22:Codebook mapping $S(\mathbf{M})$: $u^{\text{P}}_{k}=u^{\text{P}}_{k}$, $v=v+1024$, $u^{\text{O}}_{k}=u^{\text{O}}_{k}+1024+2048$, where $1024$ is the codebook size of $u^{\text{P}}_{k}$ and $2048$ is the codebook size of $v$.

### v5

@@ 3.2 Dual-Stream Point Conditioner
Following [6, 5, 7], our default input condition for Nautilus is point cloud, given its easy accessibility and rich geometric information. We design a dual-stream point conditioner that provides both global and local geometric guidance.
@@
Global Point Cloud Encoder. In order to capture high-quality global features from the input point cloud, we employ the Michelangelo point cloud encoder [50] to extract complete geometric information. As illustrated in Figure 5, the point cloud of the entire mesh is encoded into a global latent feature, $f_{\text{glb}}$, which serves as the key and value in the cross-attention layer of the decoder. This global feature encapsulates the overall geometry of the mesh asset, enabling the decoder to grasp the general shape of the input point cloud and autoregressively generate an approximate mesh output that reflects the underlying structure.
@@
Local Point Cloud Encoder. As discussed in Sec. 1, fine-grained local geometric fidelity is crucial for preserving mesh topology and structural coherence. To this end, we propose a local-aware feature extraction mechanism that aligns with our shell-based generation paradigm, providing precise geometric guidance at the vertex-fan level. As depicted in Figure 5, we employ a PointConv [47] module $f_{\text{loc}}(\cdot)$ to capture local geometric information from the point cloud input. Moreover, we design a feature injection scheme that augments token embeddings with their local geometric context, inspired by ControlAR [19]. Specifically, for each vertex $P_{k,i}$ in the $k$-th shell, we incorporate the local feature $f_{\text{loc}}(O_{k})$ extracted from its central vertex $O_{k}$ into its tokens feature. For each shell’s center vertex, we construct a local neighborhood through KNN sampling of the 100 nearest points, ensuring dense coverage of local surface geometry. This tight integration with our shell-based tokenization enables progressive incorporation of local geometric constraints during generation, enhancing both topological consistency and structural fidelity.
@@
Figure 5: Illustration of our Dual-Stream Point Conditioner. The Global Encoder extracts global features $f_{\text{glb}}$ from the entire point cloud for cross-attention layers. The Local Encoder captures fine-grained topology information via point convolution, adding local geometric features $f_{\text{loc}}$ to each token during generation.
3.3 Autoregressive Sequence Decoder
@@
The core generator of our Nautilus is a transformer decoder that autoregressively predicts the probability of the next mesh token based on the preceding.
@@
Autoregressive Mesh Token Generation. After tokenization, the mesh generation task is simply reformulated as an autoregressive sequence prediction problem [37, 31, 3]. Using the next-token prediction paradigm, models are trained to predict each discrete coordinate $c\in\{u,v\}$ based on the conditional probability given previously generated tokens. The probability of a mesh $\mathbf{M}$ can be expressed as:
$p(\mathbf{M})=\prod_{i=1}^{L}p(c_{i}|c_{<i}),\quad c_{i}\in\{0,1,\dots,\alpha-1\}$ (5)
@@
$L_{\text{CE}}=\text{CrossEntropy}(\hat{S},S(\mathbf{M})_{>0})$ (6)
@@
3.4 Image Conditioned Generation
@@
To further expand practical applications, we extend our Nautilus-style tokenization and autoregressive framework to support image-conditioned [33] mesh generation. We achieve this by leveraging the Michelangelo [50], a pre-trained shape generation framework that aligns multiple modalities in a unified feature space. Specifically, we replace the original decoder in Michelangelo with our mesh decoder, which incorporates Nautilus-style tokenization. The decoder is trained while keeping the point cloud encoder frozen, enabling mesh generation from the unified feature space. After training, we integrate Michelangelo’s image encoder to enable single-image conditioning. Using the aligned feature space, our decoder generates high-quality meshes directly from images, as demonstrated by the experimental results in Sec. 4.
4 Experiments
@@
Figure 6: Qualitative comparison on point cloud conditioned generation between Nautilus and latest open-source methods [6, 7].
@@
Figure 7: Qualitative comparison on single image conditioned generation between Nautilus and EdgeRunner [33]. The results are presented with corresponding normal maps to highlight topological details.
4.1 Implementation Details
[…]
4.2 Qualitative Results
Point Cloud Condition. We compare our Nautilus with the latest open-source direct mesh generation methods [6, 7] in the point cloud conditioned generation, where EdgeRunner [33] is excluded from this comparison due to its unavailablity of pretrained checkpoints and thus we provide a comparison discussion with EdgeRunner in Appendix. To construct the test set, we use open-source image-to-3D methods [35] to generate dense meshes and directly sample point clouds from them, ensuring that all test samples are entirely unseen during training. The selected samples present significant challenges, featuring thin structures, anisotropic faces, and intricate geometric details.
As shown in Figure 6, existing methods struggle with these challenging cases, exhibiting structural defects such as surface holes and missing components, while failing to capture fine topological details. In contrast, Nautilus generates aesthetically appealing, artist-like meshes that faithfully preserve the geometric details of the input. This performance gap stems from their inadequate locality modeling. Current tokenization approaches generate excessive sequence lengths, limiting these methods to training on simple meshes with fewer than 1,600 faces—insufficient for complex test cases. Furthermore, the limited local dependency compromises structural fidelity and increases manifold defects. Nautilus addresses these limitations by leveraging locality in both mesh tokenization and sequence generation, achieving superior quality even for complex topologies, with additional results presented in Appendix.
Single-view Image Condition. In Figure 7, we present a comparison of meshes generated from single-view images between our Nautilus and EdgeRunner [33], the only autoregressive method capable of directly conditioning on images  [^2] Note that our comparison is limited to EdgeRunner’s released samples, as their checkpoints are not publicly available.. As shown in Figure 7, Nautilus generates detailed, manifold meshes with sharp features that accurately match the input conditions, which highlights the superiority of our Nautilus-style tokenization, as further validated in Sec. 4.3. We present more image-conditioned results and a comparison discussion with EdgeRunner in Appendix.
Metrics AMT [7] EdgeRunner [33] Ours
@@ 4.2 Qualitative Results
Table 1: Quantitative comparison with other tokenization algorithms. Our Nautilus-style tokenization achieves more efficient compression while better preserving the local dependency.
@@
Table 2: Quantitative Comparison with latest open-source methods, where our Nautilus framework significantly outperform others in both condition fidelity and visual quality.
4.3 Quantitative Results
Tokenization Algorithm. In Table 1, we compare our Nautilus-style mesh tokenization algorithm with Adjacent Mesh Tokenization (AMT) [7] and EdgeRunner [33] in two aspects: (1) Compression Ratio measures the reduction in sequence length relative to the original representation $9N$. Higher compression ratios enable modeling of meshes with more faces and better topological complexity. (2) Local Ratio, a new metric that we introduce, quantifies structural coherence among neighboring tokens. For each vertex, we calculate the proportion of its neighboring vertices in the mesh that appear within 100 preceding positions in the sequence. Higher local ratios indicate better preservation of local dependencies after serialization. In Table 1, our tokenization significantly outperforms other methods in both metrics, which is attributed to the explicit locality modeling in its design. In contrast, AMT processes mesh faces in an order of elongated strips rather than interconnected local parts, neglecting local dependencies during tokenization, while EdgeRunner relies on the excessive introduction of specialized tokens, leading to suboptimal performance.
@@ 4.3 Quantitative Results
Figure 8: We conduct ablation comparison on four specific configurations: using only the Coordinate Compression (Comp. only), using only the Nautilus Shell representation (Rep. only), using our Full Tokenization but without the local point encoder, and our Full Version.
Generation Performance. We compared our point cloud conditioned generation results with MeshAnything [6] and MeshAnythingV2 [7], on a high-quality test set of 500 samples collected using the same protocol as in our qualitative experiments. Following the settings of previous work [33], we evaluate overall reconstruction quality by uniformly sampling 1,024 points from the surfaces of both ground truth and predicted meshes, and computing the Chamfer Distance (C.Dist.) and Hausdorff Distance (H.Dist.) between the sampled point clouds. Lower distances between point sets indicate higher reconstruction accuracy and better fidelity to the input condition. In addition, we conducted a user study to assess the proportion of results that users find satisfactory, with further details provided in Appendix. In Table 2, Nautilus significantly outperforms previous approaches, validating its effectiveness that is consistent with the qualitative analysis in Sec. 4.2. We present further quantitative evaluation on topological fidelity in Appendix.
4.4 Ablation and Discussion
@@
To validate the effectiveness of our proposed Nautilus-style tokenization and the Dual-Stream Point Conditioner, we perform extensive ablation studies on the tokenization scheme and the local encoder in our conditioner, where a quantitative ablation comparison is presented in Appendix.
@@
Tokenization Scheme. We evaluated two configurations to demonstrate the key components of our tokenization design: using only coordinate compression (Comp. only) and using only the Nautilus Shell representation (Rep. only). Figure 8 compares these variants with our full version. Comp. only achieves a compression ratio of 0.549 and Rep. only achieves 0.484, both substantially higher than our full version’s 0.275. In terms of generation quality, Comp. only fails to construct the topology completely, while Rep. only produces a coarse outline of the target shape. In contrast, with the complete tokenization, our full version produces precisely structured results that are consistent with the input condition in geometry. With the explicit modeling of local neighboring vertices used in our Nautilus Shell representation, both Rep. only and our full version demonstrate better performance compared to Comp. only. In addition, as shown in the comparison among the two ablated groups and our full version, insufficient compression excludes complex training samples, resulting in the model’s inability to learn intricate topologies and leading to generation failures.
@@
Local Point Encoder. To evaluate the local encoder’s effectiveness in the point conditioner, we compare our full version with a variant that uses full tokenization but excludes the local encoder (Full Tokenization). As shown in Figure 8, the variant without the local encoder exhibits minor manifold defects in complex topological regions, particularly surface holes in local structures. In contrast, our full version, which incorporates the local encoder, successfully resolves these issues and generates precise, compact meshes. The results demonstrate that the local geometric modeling provided by our local encoder yields significant improvements in local details and connectivity, underscoring the importance of enhanced local dependency for generating precise local structures in challenging regions.
5 Conclusion
@@
In this paper, we present Nautilus, a locality-aware autoencoder designed for scalable high-fidelity mesh generation. By leveraging the locality of manifold meshes, Nautilus introduces a novel tokenization algorithm that preserves local dependency while achieving significant sequence compression, as well as a novel point conditioning mechanism that enhances both global consistency and local structure fidelity. Our extensive experiments demonstrate that Nautilus outperforms state-of-the-art methods across various metrics, enabling the generation of meshes with unprecedented topological complexity and details. These results validate the importance of incorporating locality in tokenization and sequence prediction to achieve superior mesh generation.
Acknowledgments
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-030).
Supplementary Material
The Appendix is organized as follows:
  • Section A: gives a step-by-step demonstartion of our Nautilus-style Tokenization algorithm.
  • Section B: further provides additional experimental results, comparisons, and analyses, including extended visualizations, quantitative evaluations of topological quality, and more in-depth quantitative ablation studies.
  • Section C: provides in-depth discussions on Nautilus, specifically detailing the methodology and experimental comparisons with EdgeRunner.
  • Section D: presents the statistical analysis conducted on our training dataset.
  • Section E: elaborates more details about our user study.
  • Section F: discusses the limitation of our approach.
We also include a demonstration video in our supplementary materials for better visualization.
@@ Acknowledgments
Figure 9: We release additional point cloud conditioned generation results of our Nautilus across an extensive range of object categories.
@@
Figure 10: We release additional image conditioned generation results of our Nautilus across an extensive range of object categories. The results are presented alongside corresponding normal maps to highlight the topology.
Appendix A Tokenization: Step-by-Step Demonstration
@@
1:Input: manifold mesh asset $\mathcal{M}=(\mathcal{V},\mathcal{F})$ consisting of vertices $\mathcal{V}=\{v_{i}\}$ and faces $\mathcal{F}=\{f_{i}\}$.
@@
11:    Extend $S(\mathcal{M})\leftarrow S(\mathcal{M})\cup\{O_{k},P_{k,1},\dots,P_{k,\text{end}}\}$, shell $k$ complete.
@@
14:    If $\deg(N^{\text{max}}_{k,\text{end}})>4$ (ensure at least 3 connected faces in the next shell):
@@
22:Codebook mapping $S(\mathbf{M})$: $u^{\text{P}}_{k}=u^{\text{P}}_{k}$, $v=v+1024$, $u^{\text{O}}_{k}=u^{\text{O}}_{k}+1024+2048$, where 10241024 is the codebook size of $u^{\text{P}}_{k}$ and $2048$ is the codebook size of $v$.